OpenStudy (anonymous):

Hey all, I am working on #3g of 5A Problem Set 2 and I am having trouble understanding how they determined the derivative of the inverse trig function arctan(x/sqrt(1-x^2)). By definition, d/dx(arctan(x))=1/(1+x^2). Also, from the soln., I see that they took all the junk in the parentheses and called all of that "y," but isn't this function y=arctan(x/(sqrt(1-x^2))) to begin with, making it also tan(y)=x/(sqrt(1-x^2))?

OpenStudy (phi):

First, you ask, *** isn't this function y=arctan(x/(sqrt(1-x^2))) to begin with? *** Short answer: no. Longer answer: it is a convention to write y= f(x), and so you might think y = arctan( stuff) But in this case, they are using "y" as just another variable with no special meaning other than to say \[ y= \frac{x}{\sqrt{1+x^2}}\] which they do so they can make the write-up easier to read. To avoid this confusion, we can use "u" instead. Let \[ u= \frac{x}{\sqrt{1+x^2}}\] By the "chain rule" \[ \frac{d}{dx} \tan^{-1}u = \frac{d}{du} \tan^{-1}u \cdot \frac{d}{dx} u \]

OpenStudy (phi):

We get \[ \frac{d}{dx} \tan^{-1}u = \frac{d}{du} \tan^{-1}u \cdot \frac{d}{dx} u \\ = \frac{1}{1+u^2} \cdot \frac{d}{dx} \left( \frac{x}{\sqrt{1+x^2}}\right) \] we can replace u^2 with its definition and simplify the 1/(1+u^2) we can take the derivative of the other term. simplify to get the answer. Notice that the chain rule is needed to do this problem.... and it gets a bit messy

OpenStudy (anonymous):

Okay, that makes a whole lot of sense. Thank you for your reply and sorry it took me a bit to reply:]